OpenStudy (saldis):

What are the x and y-intercepts of the line described by this equation? 2x + 4y = 12.4 Enter your answers in the boxes in decimal form. x-intercept = y-intercept =

OpenStudy (mathmale):

I'm sure you've calculated x- and y-intercepts before. What did you do to obtain them?

OpenStudy (roguehamster2073):

ill let mathmale answer he was here first

OpenStudy (roguehamster2073):

unless u give me permission @mathmale

OpenStudy (saldis):

lol idk

OpenStudy (roguehamster2073):

Well

OpenStudy (roguehamster2073):

Ill answer

OpenStudy (saldis):

ok

OpenStudy (roguehamster2073):

You have to graph 2x+4y =12.4 first, before you can even get close to answering

OpenStudy (roguehamster2073):

on a coordinate graph, do you know how to do that

OpenStudy (saldis):

I am a blonde.... do you think I know how to do that

OpenStudy (roguehamster2073):

https://www.desmos.com/calculator use this

OpenStudy (roguehamster2073):

type 2x+4y =12.4 in

OpenStudy (roguehamster2073):

tell me when you are done

OpenStudy (saldis):

I am not lol

OpenStudy (roguehamster2073):

Well, if you go to the link I g ave, and type 2x+4y =12.4 in, its real simple

OpenStudy (roguehamster2073):

or copy/paste it in

OpenStudy (saldis):

Ok i did and idk what the line is ahahah

OpenStudy (mathmale):

Of course, graphing will help. However, I'd suggest an easier approach: To find the y-intercept, let x=0. To find the x-intercept, let y=0. @ saldis: Try this approach now. 2x + 4y = 12.4

OpenStudy (roguehamster2073):

Ok grea let me explain

OpenStudy (roguehamster2073):

yes

OpenStudy (roguehamster2073):

try his first before i explain

OpenStudy (mathmale):

I'd suggest you learn to do this without the help of computer software. Check your results with such software, if you wish.

OpenStudy (roguehamster2073):

Listen to him hes expert, but leave the link i gave you open

OpenStudy (mathmale):

2x + 4y = 12.4 Let x=0; find y now, please. @saldis?

OpenStudy (saldis):

ok

OpenStudy (roguehamster2073):

i dont think she knows what the graph is

OpenStudy (roguehamster2073):

or x y

OpenStudy (saldis):

Yea your right I am a blonde I dont know what colors are either XD

OpenStudy (roguehamster2073):

Ok saldis do you still have the link I gave you with the line open

OpenStudy (roguehamster2073):

do you?

OpenStudy (saldis):

no I got the answer thank you

OpenStudy (roguehamster2073):

Great

OpenStudy (mathmale):

Still, I think you'd better learn how to find x- and y-intercepts quickly and accurately; you will need this skill in future problems.